OpenStudy (anonymous):

what is the solution to this equation? 3+1/4x=2

OpenStudy (sashi786):

0.5

OpenStudy (anonymous):

1/4 = .25 3 + .25x = 2 \[.25x + 3 = 2\] \[.25x + 3 - 3 = 2 - 3\] \[.25x = -1\] \[\frac{ .25 }{ .25 } = \frac{ -1 }{ .25 }\] \[x = -4\]

OpenStudy (anonymous):

if this was adequit please give me a medal if you dont mind